Advertisements
Advertisements
प्रश्न
A card from a pack of 52 playing cards is lost. From the remaining cards of the pack three cards are drawn at random (without replacement) and are found to be all spades. Find the probability of the lost card being a spade.
Advertisements
उत्तर
Let E1, E2, E3, E4 and A be the event defined as below:
E1 = the missing card is a heart card
E2 = the missing card is a spade card
E3 = the missing card is a club card
E4 = the missing card is a diamond card
A = drawing three spade cards from the remaining cards
Now, we have the following:
`P(E_1)=13/52=1/4`
`P(E_2)=13/52=1/4`
`P(E_3)=13/52=1/4`
`P(E_4)=13/52=1/4`
`P(A/E_1)=(""^13C_3)/(""^51C_3)`
`P(A/E_2)=(""^12C_3)/(""^51C_3)`
`P(A/E_3)=(""^13C_3)/(""^51C_3)`
`P(A/E_4)=(""^13C_3)/(""^51C_3)`
By Bayes Theorem, we have:
Required probability =`P(E_2/A)`
`=(P(E_1)P(A/E_2))/(P(A/E_1)E_1+P(A/E_2)E_2+P(A/E_3)E_3+P(A/E_4)E_4)`
`=(1/4xx(""^13C_3)/(""^51C_3))/((""^13C_3)/(""^51C_3)xx1/4+(""^12C_3)/(""^51C_3)xx1/4+(""^13C_3)/(""^51C_3)xx1/4+(""^13C_3)/(""^51C_3)xx1/4)`
`=(""^12C_3)/(""^12C_3+3xx""^13C_3)`
`=220/(220+286xx3)=110/539`
APPEARS IN
संबंधित प्रश्न
A bag contains 4 balls. Two balls are drawn at random (without replacement) and are found to be white. What is the probability that all balls in the bag are white?
A die, whose faces are marked 1, 2, 3 in red and 4, 5, 6 in green is tossed. Let A be the event "number obtained is even" and B be the event "number obtained is red". Find if A and B are independent events.
One card is drawn at random from a well-shuffled deck of 52 cards. In which of the following case is the events E and F independent?
E : ‘the card drawn is a king or queen’
F : ‘the card drawn is a queen or jack’
Two events, A and B, will be independent if ______.
Prove that if E and F are independent events, then the events E and F' are also independent.
A speaks the truth in 60% of the cases, while B is 40% of the cases. In what percent of cases are they likely to contradict each other in stating the same fact?
A problem in statistics is given to three students A, B, and C. Their chances of solving the problem are `1/3`, `1/4`, and `1/5` respectively. If all of them try independently, what is the probability that, problem is solved?
A, B, and C try to hit a target simultaneously but independently. Their respective probabilities of hitting the target are `3/4, 1/2` and `5/8`. Find the probability that the target
- is hit exactly by one of them
- is not hit by any one of them
- is hit
- is exactly hit by two of them
The probability that a student X solves a problem in dynamics is `2/5` and the probability that student Y solves the same problem is `1/4`. What is the probability that
- the problem is not solved
- the problem is solved
- the problem is solved exactly by one of them
A bag contains 3 yellow and 5 brown balls. Another bag contains 4 yellow and 6 brown balls. If one ball is drawn from each bag, what is the probability that, both the balls are of the same color?
Bag A contains 3 red and 2 white balls and bag B contains 2 red and 5 white balls. A bag is selected at random, a ball is drawn and put into the other bag, and then a ball is drawn from that bag. Find the probability that both the balls drawn are of same color
The probability of simultaneous occurrence of at least one of two events A and B is p. If the probability that exactly one of A, B occurs is q, then prove that P(A′) + P(B′) = 2 – 2p + q.
Two dice are thrown together. Let A be the event ‘getting 6 on the first die’ and B be the event ‘getting 2 on the second die’. Are the events A and B independent?
If A and B′ are independent events then P(A′ ∪ B) = 1 – ______.
Refer to Question 1 above. If the die were fair, determine whether or not the events A and B are independent.
Let E1 and E2 be two independent events such that P(E1) = P1 and P(E2) = P2. Describe in words of the events whose probabilities are: (1 – P1) P2
Let E1 and E2 be two independent events such that P(E1) = P1 and P(E2) = P2. Describe in words of the events whose probabilities are: 1 – (1 – P1)(1 – P2)
In Question 64 above, P(B|A′) is equal to ______.
If A and B are independent, then P(exactly one of A, B occurs) = P(A)P(B') + P(B)P(A')
One card is drawn at random from a well-shuffled deck of 52 cards. In which of the following case is the events E and F independent?
E : ‘the card drawn is black’
F : ‘the card drawn is a king’
Given two independent events A and B such that P(A) = 0.3, P(B) = 0.6 and P(A' ∩ B') is ______.
Events A and Bare such that P(A) = `1/2`, P(B) = `7/12` and `P(barA ∪ barB) = 1/4`. Find whether the events A and B are independent or not.
Two players A and B are alternately throwing a coin and a die together. A player who first throws head and 6 wins the game. If A starts the game, then the probability that B wins the game is ______.
What are independent events in probability?
For independent events A and B, what is the probability that at least one occurs?
A die is thrown once. What is the sample space \[S\]?
For \[E=\{3,6\}\] and \[F=\{2,4,6\}\], what is \[E\cap F\]?
Given \[P(E\cap F)=\frac{1}{6}\] and \[P(E)P(F)=\frac{1}{6}\], what is the relation between E and F?
